Mathematics 242 – Principles of Analysis Definitions Quiz 4 – October 14, 2005

Your Name:

1. What is the definition of convergence for a sequence (x_{n})?

2. Give an example of a sequence containing one subsequence that converges and a different subsequence that diverges.